# Multiscale Dynamical Indices Reveal Scale-Dependent Atmospheric Dynamics

### v1

@@ Acknowledgments
In this section of the supplementary text, we expand on the definition of instantaneous dimension ($d$) and inverse persistence ($\theta$) used in the main text. As introduced in the main text, these two dynamical indices grounded at the intersection of dynamical systems theory and extreme value theory, and were developed by [12] and [13] to characterize local dynamical properties of complex systems. For one given state of interest $\boldsymbol{\zeta}$ in the phase space, this methodology uses its neighboring states, also known as recurrences, to derive its state-dependent dynamical properties. Specifically, for the considered state $\boldsymbol{\zeta}$, we first calculate its negative logarithmic distances to other states:
@@
where ${\bf{x}}(t)$ is the trajectory of the studied system and the $dist$ function can be any distance metric. Requiring a trajectory falls within a neighborhood of $\boldsymbol{\zeta}$ (a hypersphere in the phase space) is a synonym of having the time series $g[{\boldsymbol{\zeta}},{\bf{x}}(t)]$ above one threshold $s(q,{\boldsymbol{\zeta}})$, where $s(q,{\boldsymbol{\zeta}})$ is a high threshold associated with a quantile $q$ of the series ${\bf{X}}\equiv g({\boldsymbol{\zeta}},{\bf{x}}(t))$ ($q$ is chosen as a high value to identify neighboring states). Then, we define a quantity called exceedance for the neighboring states (i.e., recurrences) of $\boldsymbol{\zeta}$ as follows:
@@
For the other local index, inverse persistence $\theta$, its definition is adopted from the extremal index in extreme value theory. This dimensionless parameter quantifies the inverse of the clustering duration of extremes. In this context, extremes are defined as recurrences within the neighborhood of the reference state ${\boldsymbol{\zeta}}$. Consequently, persistence ($\theta^{-1}$) represents the average number of consecutive recurrences, thereby providing insight into the system’s instantaneous persistence. Unlike the local dimension ($d$), which requires exceedances for its estimation, the inverse persistence ($\theta$) is derived solely from the chronological order of recurrences of the reference state ${\boldsymbol{\zeta}}$. The full mathematical derivation of these two indices can be found in [33].
Text S2. Definition of European Regional Heatwaves
We define regional warm-temperature extremes over the British Isles, Scandinavia, central Europe (Germany), Iberia, and the central Mediterranean for extended summer season (JJAS), following the bounding boxes used in [28].
To this end, we first define two criteria for each individual grid point using ERA5 daily maximum 2-meter temperature data from 1979 to 2022, following [29]. The first criterion requires the daily maximum temperature to exceed a threshold, defined as the 90th percentile of the daily maximum temperature for that calendar day, calculated using a centered 15-day window. The second criterion requires $M_{d}>0$, where $M_{d}$ is a daily heatwave magnitude index defined as:
@@
Text S3. Definition of Atmospheric Blocking events
@@
We applied a standard blocking detection algorithm based on the reversal of 500-hPa geopotential height gradients, as described by [56] and references therein. For each grid point, we first compute its geopotential height gradients to the north ($\Delta Z_{\mathrm{N}}$) and to the south ($\Delta Z_{\mathrm{S}}$):
@@
where $Z$ represents geopotential height at 500-hPa, $\Delta\phi=15^{\circ}$, and $\phi$ ranges from $50^{\circ}N$ to $75^{\circ}N$. Instantaneous blocking (IB) is defined at a specific grid point if the gradients simultaneously satisfy the following conditions:
@@
Next, we apply spatiotemporal filtering to the IB field to extract large-scale and slow-moving events. The maximum IB index within ±$4^{\circ}$ latitude is taken to account for meridional movement. Extended IB cases are then selected if they span at least $15^{\circ}$ longitude, filtering out systems that are too small. Finally, blocking is defined as an extended IB persisting within ±$10^{\circ}$ longitude for at least five consecutive days, ensuring the detection of only persistent and slow-moving systems.
Blocked days are further determined for three 30° longitude regions, namely Greenland ($60^{\circ}$ to $30^{\circ}$), North Atlantic ($30^{\circ}$ to $0^{\circ}$), and Scandinavian ($0^{\circ}$ to $30^{\circ}$). A day is considered blocked in a given region if a block spans more than half of the region (i.e., exceeds 15° of longitude within the region). Based on this, we can plot the composite anomalies of the standardized anomalies for blocked days within the three regions, as shown in Fig. 10.
Figure 7: Temporal evolution of the composite anomaly of Z500 before the onset and after the decay of European summertime heatwaves. Composite anomaly of Z500 from three days before the onset to the onset day, and from the last day to three days after, of summertime heatwaves across Europe.
Figure 8: Temporal evolution of the composite anomaly of Z500-based local dimension ($d$) before the onset and after the decay of European summertime heatwaves. As in Fig. S3, but for Z500-based local dimension.
Figure 9: Temporal evolution of the composite anomaly of Z500-based inverse persistence ($\theta$) before the onset and after the decay of European summertime heatwaves. As in Fig. S3, but for Z500-based inverse persistence.

### v2

@@ Appendix A Dynamical Indices
In this section of the supplementary text, we expand on the definition of instantaneous dimension ($d$) and inverse persistence ($\theta$) used in the main text. As introduced in the main text, these two dynamical indices grounded at the intersection of dynamical systems theory and extreme value theory, and were developed by Lucarini et al. [7] and Faranda et al. [8] to characterize local dynamical properties of complex systems. For one given state of interest $\boldsymbol{\zeta}$ in the phase space, this methodology uses its neighboring states, also known as recurrences, to derive its state-dependent dynamical properties. Specifically, for the considered state $\boldsymbol{\zeta}$, we first calculate its negative logarithmic distances to other states:
@@
where ${\bf{x}}(t)$ is the trajectory of the studied system and the distdist function can be any distance metric. Requiring a trajectory falls within a neighborhood of $\boldsymbol{\zeta}$ (a hypersphere in the phase space) is a synonym of having the time series $g[{\boldsymbol{\zeta}},{\bf{x}}(t)]$ above one threshold $s(q,{\boldsymbol{\zeta}})$, where $s(q,{\boldsymbol{\zeta}})$ is a high threshold associated with a quantile $q$ of the series ${\bf{X}}\equiv g({\boldsymbol{\zeta}},{\bf{x}}(t))$ ($q$ is chosen as a high value to identify neighboring states). Then, we define a quantity called exceedance for the neighboring states (i.e., recurrences) of $\boldsymbol{\zeta}$ as follows:
@@
For the other local index, inverse persistence $\theta$, its definition is adopted from the extremal index in extreme value theory. This dimensionless parameter quantifies the inverse of the clustering duration of extremes. In this context, extremes are defined as recurrences within the neighborhood of the reference state $\boldsymbol{\zeta}$}. Consequently, persistence ($\theta^{-1}$) represents the average number of consecutive recurrences, thereby providing insight into the system’s instantaneous persistence. Unlike the local dimension ($d$), which requires exceedances for its estimation, the inverse persistence ($\theta$) is derived solely from the chronological order of recurrences of the reference state $\boldsymbol{\zeta}$}. The full mathematical derivation of these two indices can be found in Faranda et al. [28].
Appendix B Definition of European Regional Heatwaves
We define regional warm-temperature extremes over the British Isles, Scandinavia, central Europe (Germany), Iberia, and the central Mediterranean for extended summer season (JJAS), following the bounding boxes used in Zschenderlein et al. [23].
To this end, we first define two criteria for each individual grid point using ERA5 daily maximum 2-meter temperature data from 1979 to 2022 [24]. The first criterion requires the daily maximum temperature to exceed a threshold, defined as the 90th percentile of the daily maximum temperature for that calendar day, calculated using a centered 15-day window. The second criterion requires $M_{d}>0$, where $M_{d}$ is a daily heatwave magnitude index defined as:
@@ Appendix B Definition of European Regional Heatwaves
Appendix C Definition of Atmospheric Blocking events
@@
We applied a standard blocking detection algorithm based on the reversal of 500 hPa geopotential height gradients, as described by Brunner et al. [53] and references therein. For each grid point, we first compute its geopotential height gradients to the north ($\Delta Z_{\mathrm{N}}$) and to the south ($\Delta Z_{\mathrm{S}}$):
@@
where $Z$ represents geopotential height at 500 hPa, $\Delta\phi=15^{\circ}$, and $\phi$ ranges from $50^{\circ}N$ to $75^{\circ}N$. Instantaneous blocking (IB) is defined at a specific grid point if the gradients simultaneously satisfy the following conditions:
@@
Next, we apply spatiotemporal filtering to the IB field to extract large-scale and slow-moving events. The maximum IB index within $4^{\circ}$ latitude is taken to account for meridional movement. Extended IB cases are then selected if they span at least $15^{\circ}$ longitude, filtering out systems that are too small. Finally, blocking is defined as an extended IB persisting within $10^{\circ}$ longitude for at least five consecutive days, ensuring the detection of only persistent and slow-moving systems.
Blocked days are further determined for three 30° longitude regions, namely Greenland ($60^{\circ}$W to $30^{\circ}$W), North Atlantic ($30^{\circ}$W to $0^{\circ}$), and Scandinavian ($0^{\circ}$ to $30^{\circ}$E). A day is considered blocked in a given region if a block spans more than half of the region (i.e., exceeds 15° of longitude within the region). Based on this, we can plot the composite anomalies of the standardized anomalies for blocked days within the three regions, as shown in Fig. 9.